OpenStudy (anonymous):

5.The surface area of a sphere is 324pi m². What is the volume of the sphere

OpenStudy (anonymous):

find the radius and then use it to find the volume. this is similar to the problem you just posted. find the radius in the same way and then use it in the volume formula for a sphere.

OpenStudy (anonymous):

\[V_{Sphere} = \frac{ 4 }{ 3 }\pi r^3\]

OpenStudy (anonymous):

wait so wat would sphere in this problem be

OpenStudy (anonymous):

324pi m= 4pi r^2 divide 324 by 4 you get 81 then find the square root of 81 which is 9 that would be your radius then plug in 9 as your radius for the volume of the sphere. V=(4/3)pi times 9 to the third power give you V=(4/3)pi x 729 divide by (4/3) give you 546.75pi would be your answer is that rite

OpenStudy (agent0smith):

\[\Large SA = 4 \pi r^2\]first find the radius, your SA is 324

OpenStudy (anonymous):

yeah that looks great except you multiply by 4/3 not divide.

OpenStudy (anonymous):

so wat would the answer be if i multiplied

OpenStudy (anonymous):

come on man...

OpenStudy (anonymous):

\[\frac{ 4 }{ 3 }\cdot 729\cdot \pi\]

OpenStudy (anonymous):

@futureCollegeStudent use a calculator:)

OpenStudy (anonymous):

You got this!!!!!!!!!!

OpenStudy (anonymous):

\[\frac{ 4 }{\cancel{3} }\cdot \frac{ 9\cdot9\cdot(3\cdot\cancel{3})\cdot \pi }{ 1 }=\]

OpenStudy (anonymous):

3052.08

OpenStudy (anonymous):

close but no cigar... did you round pi? \[=972\,\pi\]

OpenStudy (anonymous):

i get the same answer every time :(

OpenStudy (anonymous):

what do you use for pi?

OpenStudy (anonymous):

you're using 3.14 for pi... use pi as stored in your calculator or learn a fractional equivalent or more deciamla places in your approximation. try using 3.141592654

OpenStudy (anonymous):

"decimal"